# Is the MMI Criterion Necessary for Interpretability? Degenerating Non-causal Features to Plain Noise for Self-Rationalization

### v1

@@ 1 Introduction
One important line of research to build self-explainable NLP models is first extracting the most informative rationale in a text and then using the extracted rationale to train a predictor. This line of research is known as rationalization. A model-agnostic rationalization framework, called Rationalizing Neural Predictions (RNP), was first proposed by Lei et al. (2016). RNP utilizes a cooperative game between an extractor and a predictor. This game is designed with a focus on "data-centric" importance of rationales (i.e., it aims to explain the connection between a text and the model-agnostic task label, rather than explaining the output of a specific model). First, the extractor identifies the most informative part of the input, known as the rationale. Then, as depicted in Figure 1, the rationale is transmitted to the predictor to facilitate predictions. The extractor and predictor are trained cooperatively to maximize prediction accuracy, with the theoretical support being the Maximum Mutual Information (MMI) criterion (Yu et al., 2021; Chang et al., 2020). RNP and its variants have become mainstream approaches for enhancing the interpretability of NLP models (Yue et al., 2023; Liu et al., 2023b; Storek et al., 2023; Zhang et al., 2023). Aside from interpretability, rationalization can also serve as a method for data cleaning, as the extracted $(Z,Y)$ samples can function as a new dataset. Recent studies have shown that a predictor trained with such a dataset can be more robust (Chen et al., 2022) and generalizable (Wu et al., 2022; Gui et al., 2023), due to the removal of task-irrelevant, harmful information.
@@
Some methods try to develop regularizers that can penalize the spurious features and fix the shortcoming of MMI. INVRAT (Chang et al., 2020) incorporates the concept of invariant risk minimization to design an invariance penalty. Inter_RAT (Yue et al., 2023) utilizes an intervention penalty. CR (Zhang et al., 2023) implements a sufficiency and necessity penalty by separately assessing the sufficiency and necessity of each token. In addition to the specific shortcomings of each type of method (discussed in §$\S$2), they share a common limitation: most still adhere to the MMI criterion and merely use supplementary objectives to penalize spurious features. If the penalty term’s weight is too small, spurious features will still be favored over uninformative noise due to their higher mutual information. Consequently, when the extractor initially selects noise, the gradient descent algorithm might shift towards either spurious features or the true rationale. On the other hand, if the penalty term’s weight is too high, it can dominate the loss function and impair the MMI’s ability to distinguish between noise and causal features (see §$\S$4.1). The difference between spurious features and noise can complicate the loss landscape of rationale extraction, which may lead to the emergence of local optima. Note that the problem of local optima in rationalization is very serious (Yu et al., 2021). A recent research MCD (Liu et al., 2023a) revises the MMI criterion to the minimum conditional dependence criterion and does not introduce extra penalty terms. However, MCD also can not promise to treat the spurious features as plain noise. We provide a detailed comparison of MCD and our approach in Appendix A.1 help readers better understand the unique advantages of our approach.
@@ 2 Related work
Mitigating spurious correlations. One important obstacle of rationalization is the spurious features in datasets, as the spurious features also have high correlations with the task label and can compete with the causal features for extraction opportunities under the most widely used MMI criterion. Some methods have been developed to mitigate the impact of spurious correlations. INVRAT (Chang et al., 2020) attempts to tackle feature correlation using invariant risk minimization (IRM) (Arjovsky et al., 2019). The main idea is to penalize spurious (non-causal) variations by splitting the dataset into distinct environments. However, IRM-based methods have several limitations. For instance, they require strong prior knowledge about the relationships between non-causal and causal features (e.g., the extra labels of non-causal features) in order to divide the dataset (Lin et al., 2022b). Moreover, IRM-based methods are limited to addressing only a finite set of predetermined non-causal features, neglecting the potential existence of numerous unknown non-causal features. In fact, a recent study (Lin et al., 2022b) in the field of IRM has theoretically demonstrated that it is nearly impossible to partition a dataset into different environments to eliminate all non-causal features using IRM. Other challenges, such as the tendency to overfit, difficulty in applying to larger models (Zhou et al., 2022; Lin et al., 2022a), and the marginal shift risk of the input (Rosenfeld et al., 2021), have also been identified within the realm of IRM. Inter_RAT (Yue et al., 2023) attempts to eliminate feature correlation through backdoor adjustment, intervening directly with the confounders. However, it is extremely hard to measure the confounders since they are usually not observable in the dataset. CR (Zhang et al., 2023) calculates the sufficiency and necessity of each token separately, which leads to a high computational complexity, making it feasible only for very short texts. Aside from the above shortcomings, penalty-based methods share a common limitation. They need to coordinate the MMI and the penalty objectives to make the gradient descent algorithm treat the spurious features and plain noise equally and guide the extractor to move towards only the causal features. A recent research MCD (Liu et al., 2023a) revises MMI to the minimum conditional dependence criterion. Although MCD does not involve penalty regularizers, it also cannot treat spurious features and plain noise equally. And the spurious features can still compete with the causal ones.
@@ 3.2 Causality
We note that the contribution of this part does not belong to this paper. To help readers unfamiliar with causality better understand the spurious correlations, we borrow it from a previous paper MCD (Liu et al., 2023a) and make some minor revisions to make this paper self-contained. We provide a detailed comparison with MCD in Appendix A.1.
@@ 4.2 Spurious features are equivalent to plain noise in a counterfactual view
We denote the non-causal subset of $X$ as $A=\{S,N\}$. From the probabilistic graphical model shown in Figure 2(a), we know that $A$ and $Y$ are d-separated by the causal features $C$ (Liu et al., 2023a) (please refer to Appendix A.6 for a detailed illustration). It means that all variables within $A$ are independent with $Y$ when conditioned on $C$.
@@ 5 The practical method
The followed problem is how to apply MRD in practice. The real distributions of $P(Y|X_{-Z})$ and $P(Y|X)$ are not directly accessible. So we need further efforts to approximate them. Similar to the vanilla RNP’s approximating entropy with cross-entropy and inspired by the MCD’s (Liu et al., 2023a) success in approximating real distributions with a predictor’s output, we try to approximate real distributions by making use of the predictor. We first approximate $P(Y|X)$ with $P(\hat{Y}_{X}|X)$ by minimizing $H_{c}(Y,\hat{Y}_{X}|X)$ (please refer to Appendix A.7 for detailed analysis on the feasibility of this approximation), and we also approximate $P(Y|X_{-Z})$ with $P(\hat{Y}_{-Z}|X_{-Z})$ by minimizing the cross-entropy $H_{c}(Y,\hat{Y}_{-Z}|X_{-Z})$, where $\hat{Y}_{-Z},\hat{Y}_{X}$ are the predictor’s outputs with the inputs being $X_{-Z}$ and $X$, respectively.
@@ 6.2 Baselines and implementation details
We compare with various recent methods to show the competitiveness of our method. These methods include INVRAT (Chang et al., 2020), Inter_RAT (Yue et al., 2023), CR (Zhang et al., 2023), MCD (Liu et al., 2023a), NIR (Storek et al., 2023). Both the extractor and the predictor are composed of an encoder (e.g., RNN/Transformer) and a linear layer. We use two types of encoders: GRUs (following INVRAT, Inter_RAT, and MCD, Table 1, 2, and 3) and bert-base-uncased (following CR, Table 4). We adopt three levels of rationale sparsity: $10\%,20\%,30\%$ (achieved by adjusting $s$ in Equation 4). We report the results of five random seeds. More details are in Appendix A.9.
@@ A.1 The comparison between MCD and MRD
This paper is inspired by a previous paper MCD (Liu et al., 2023a). This part aims to clarify the distinct contributions of our MRD.
@@ A.2 A toy example of the backdoor path
This example is provided by (Liu et al., 2023a). To make the readers that are not familiar with causality better understand the spurious correlations, we borrow it to provide a more intuitive understanding of the correlation in Figure 2(b). We assume $U$, $X_{A}$, $X_{T}$, and $Y$ are all Bernoulli variables, with their respective probability distributions as:
@@ A.6 D-separation
Liu et al. (2023a) have theoretically shown that $Y$ and the non-causal features are d-separated by the causal features.
@@ A.9 More details
All datasets are in English. We process the datasets in the same way as MCD (Liu et al., 2023a). The maximum text length is set to 256. More statistics of the datasets are in Table 5. The datasets of BeerAdvocate is unbalanced. For the training data, we sample from the positive data to get same number of positive and negative texts.

### v2

@@ 1 Introduction
One important line of research to build self-explainable NLP models is first extracting the most informative rationale in a text and then using the extracted rationale to train a predictor. This line of research is known as rationalization. A model-agnostic rationalization framework, called Rationalizing Neural Predictions (RNP), was first proposed by Lei et al. (2016). RNP utilizes a cooperative game between an extractor and a predictor. This game is designed with a focus on "data-centric" importance of rationales (i.e., it aims to explain the connection between a text and the model-agnostic task label, rather than explaining the output of a specific model). First, the extractor identifies the most informative part of the input, known as the rationale. Then, as depicted in Figure 1, the rationale is transmitted to the predictor to facilitate predictions. The extractor and predictor are trained cooperatively to maximize prediction accuracy, with the theoretical support being the Maximum Mutual Information (MMI) criterion (Yu et al., 2021; Chang et al., 2020). RNP and its variants have become mainstream approaches for enhancing the interpretability of NLP models (Liu et al., 2022, 2023c, 2023a; Storek et al., 2023; Zhang et al., 2023; Liu et al., 2024; Zhao et al., 2024; Jiang et al., 2024; Hu and Yu, 2024; Yue et al., 2024). Aside from interpretability, rationalization can also serve as a method for data cleaning, as the extracted $(Z,Y)$ samples can function as a new dataset. Recent studies have shown that a predictor trained with such a dataset can be more robust (Chen et al., 2022) and generalizable (Wu et al., 2022; Gui et al., 2023), due to the removal of task-irrelevant, harmful information.
@@
Some methods try to develop regularizers that can penalize the spurious features and fix the shortcoming of MMI. INVRAT (Chang et al., 2020) incorporates the concept of invariant risk minimization to design an invariance penalty. Inter_RAT (Yue et al., 2023) utilizes an intervention penalty. CR (Zhang et al., 2023) implements a sufficiency and necessity penalty by separately assessing the sufficiency and necessity of each token. In addition to the specific shortcomings of each type of method (discussed in §$\S$2), they share a common limitation: most still adhere to the MMI criterion and merely use supplementary objectives to penalize spurious features. If the penalty term’s weight is too small, spurious features will still be favored over uninformative noise due to their higher mutual information. Consequently, when the extractor initially selects noise, the gradient descent algorithm might shift towards either spurious features or the true rationale. On the other hand, if the penalty term’s weight is too high, it can dominate the loss function and impair the MMI’s ability to distinguish between noise and causal features (see §$\S$4.1). The difference between spurious features and noise can complicate the loss landscape of rationale extraction, which may lead to the emergence of local optima. Note that the problem of local optima in rationalization is very serious (Yu et al., 2021). A recent research MCD (Liu et al., 2023b) revises the MMI criterion to the minimum conditional dependence criterion and does not introduce extra penalty terms. However, MCD also can not promise to treat the spurious features as plain noise. We provide a detailed comparison of MCD and our approach in Appendix A.1 help readers better understand the unique advantages of our approach.
@@ 2 Related work
Mitigating spurious correlations. One important obstacle of rationalization is the spurious features in datasets, as the spurious features also have high correlations with the task label and can compete with the causal features for extraction opportunities under the most widely used MMI criterion. Some methods have been developed to mitigate the impact of spurious correlations. INVRAT (Chang et al., 2020) attempts to tackle feature correlation using invariant risk minimization (IRM) (Arjovsky et al., 2019). The main idea is to penalize spurious (non-causal) variations by splitting the dataset into distinct environments. However, IRM-based methods have several limitations. For instance, they require strong prior knowledge about the relationships between non-causal and causal features (e.g., the extra labels of non-causal features) in order to divide the dataset (Lin et al., 2022b). Moreover, IRM-based methods are limited to addressing only a finite set of predetermined non-causal features, neglecting the potential existence of numerous unknown non-causal features. In fact, a recent study (Lin et al., 2022b) in the field of IRM has theoretically demonstrated that it is nearly impossible to partition a dataset into different environments to eliminate all non-causal features using IRM. Other challenges, such as the tendency to overfit, difficulty in applying to larger models (Zhou et al., 2022; Lin et al., 2022a), and the marginal shift risk of the input (Rosenfeld et al., 2021), have also been identified within the realm of IRM. Inter_RAT (Yue et al., 2023) attempts to eliminate feature correlation through backdoor adjustment, intervening directly with the confounders. However, it is extremely hard to measure the confounders since they are usually not observable in the dataset. CR (Zhang et al., 2023) calculates the sufficiency and necessity of each token separately, which leads to a high computational complexity, making it feasible only for very short texts. Aside from the above shortcomings, penalty-based methods share a common limitation. They need to coordinate the MMI and the penalty objectives to make the gradient descent algorithm treat the spurious features and plain noise equally and guide the extractor to move towards only the causal features. A recent research MCD (Liu et al., 2023b) revises MMI to the minimum conditional dependence criterion. Although MCD does not involve penalty regularizers, it also cannot treat spurious features and plain noise equally. And the spurious features can still compete with the causal ones.
@@ 3.2 Causality
We note that the contribution of this part does not belong to this paper. To help readers unfamiliar with causality better understand the spurious correlations, we borrow it from a previous paper MCD (Liu et al., 2023b) and make some minor revisions to make this paper self-contained. We provide a detailed comparison with MCD in Appendix A.1.
@@ 4.2 Spurious features are equivalent to plain noise in a counterfactual view
We denote the non-causal subset of $X$ as $A=\{S,N\}$. From the probabilistic graphical model shown in Figure 2(a), we know that $A$ and $Y$ are d-separated by the causal features $C$ (Liu et al., 2023b) (please refer to Appendix A.6 for a detailed illustration). It means that all variables within $A$ are independent with $Y$ when conditioned on $C$.
@@ 5 The practical method
The followed problem is how to apply MRD in practice. The real distributions of $P(Y|X_{-Z})$ and $P(Y|X)$ are not directly accessible. So we need further efforts to approximate them. Similar to the vanilla RNP’s approximating entropy with cross-entropy and inspired by the MCD’s (Liu et al., 2023b) success in approximating real distributions with a predictor’s output, we try to approximate real distributions by making use of the predictor. We first approximate $P(Y|X)$ with $P(\hat{Y}_{X}|X)$ by minimizing $H_{c}(Y,\hat{Y}_{X}|X)$ (please refer to Appendix A.7 for detailed analysis on the feasibility of this approximation), and we also approximate $P(Y|X_{-Z})$ with $P(\hat{Y}_{-Z}|X_{-Z})$ by minimizing the cross-entropy $H_{c}(Y,\hat{Y}_{-Z}|X_{-Z})$, where $\hat{Y}_{-Z},\hat{Y}_{X}$ are the predictor’s outputs with the inputs being $X_{-Z}$ and $X$, respectively.
@@ 6.2 Baselines and implementation details
We compare with various recent methods to show the competitiveness of our method. These methods include INVRAT (Chang et al., 2020), Inter_RAT (Yue et al., 2023), CR (Zhang et al., 2023), MCD (Liu et al., 2023b), NIR (Storek et al., 2023). Both the extractor and the predictor are composed of an encoder (e.g., RNN/Transformer) and a linear layer. We use two types of encoders: GRUs (following INVRAT, Inter_RAT, and MCD, Table 1, 2, and 3) and bert-base-uncased (following CR, Table 4). We adopt three levels of rationale sparsity: $10\%,20\%,30\%$ (achieved by adjusting $s$ in Equation 4). We report the results of five random seeds. More details are in Appendix A.9.
@@ A.1 The comparison between MCD and MRD
This paper is inspired by a previous paper MCD (Liu et al., 2023b). This part aims to clarify the distinct contributions of our MRD.
@@ A.2 A toy example of the backdoor path
This example is provided by (Liu et al., 2023b). To make the readers that are not familiar with causality better understand the spurious correlations, we borrow it to provide a more intuitive understanding of the correlation in Figure 2(b). We assume $U$, $X_{A}$, $X_{T}$, and $Y$ are all Bernoulli variables, with their respective probability distributions as:
@@ A.6 D-separation
Liu et al. (2023b) have theoretically shown that $Y$ and the non-causal features are d-separated by the causal features.
@@ A.9 More details
All datasets are in English. We process the datasets in the same way as MCD (Liu et al., 2023b). The maximum text length is set to 256. More statistics of the datasets are in Table 5. The datasets of BeerAdvocate is unbalanced. For the training data, we sample from the positive data to get same number of positive and negative texts.